# Relightable Neural Actor with Intrinsic Decomposition and Pose Control

### v1

@@ 3.3 Neural Intrinsic Decomposition
We next perform intrinsic decomposition to obtain the normal, visibility, albedo, and roughness components, required for rendering the neural actor under new light conditions. Specifically, for each point in the scene we want to obtain the normal direction $\mathbf{n}\in\mathbb{R}^{3}$, the soft visibility coefficients $\bm{\upsilon}\in\mathbb{R}^{512}$ respective to the light sources, the RGB albedo $\mathbf{a}\in\mathbb{R}^{3}$, and the scalar roughness $\rho$. The normal and visibility values can be obtained from the geometry model. However, computing the visibility of each point in 3D w.r.t. the light sources is computationally intractable for a dynamic and long sequence, since for every point in 3D and for every frame, $512$ rays have to be cast towards each pixel of the environment map. Therefore, we use a different approach.
@@
We notice that the intrinsic decomposition is only relevant for points close to the surface. Since the coarse human mesh provides a mapping from the mesh surface to UV space, and our geometry model provides a density field that models the geometry of the actor, we propose an aggregation strategy that samples normal and visibility from a given camera (therefore, only a few points are sampled in 3D) and aggregates the obtained values in a per-frame UV map. We next explain the details of this step.
3.3.1 Normal Maps
@@
For each point in a ray, we want to obtain the normal value that will be required for solving the rendering equation during relighting (will be introduced in Sec. 3.4). This value could be obtained by computing the gradient from the neural field [49]. However, this process requires back-propagating gradients throughout the entire network and often results in noisy normals. Instead, we explicitly compute the normal vector close to the visible surface by leveraging the estimate unbiased depth $\hat{\mathbf{D}}$ [43]:
$\hat{\mathbf{D}}(\mathbf{r})=\frac{\sum_{n=1}^{N}{w_{n}\delta_{n}}}{\sum_{n=1}^{N}{w_{n}}}.$ (4)
@@
Figure 3: From the obtained unbiased depth from Eq. 4, we project the depth values from camera space to 3D coordinates and approximate the tangent surface at the intersecting point, from which the normal vector is obtained. We sample multiple camera views and aggregate the depth values in UV space.
@@
Then, as illustrated in Fig. 3, we convert the depth values relative to the camera to 3D coordinates. The normal direction is obtained by computing the tangent surface at the intersection point. This process can be performed for multiple camera views to cover the surface of the human.
@@
Once the normal values are obtained in 3D, we want to project them to the UV map and aggregate them with the obtained normals from different views. In this process, the $(u^{*},v^{*})$ coordinate where the obtained normal will be projected to in the UV map is defined by
$(u^{*},v^{*})=\operatorname*{arg\,max}_{(u_{n},v_{n})}w_{n},$ (5)
@@
which corresponds to the $(u_{n},v_{n})$ value obtained from the point $\mathbf{r}_{n}$ on the ray that has the largest density weight $w_{n}$. This process ensures that the obtained normal map aggregates the normals obtained from the regions with the highest probability of being close to the surface.
@@
However, the resulting UV map can be sparse, which is undesirable for free-viewpoint rendering. Thus, the NormalNet is used to inpaint the missing normals in UV space. This lightweight CNN is implemented with $8$ layers of partial convolutions [24], takes as input sparse normal maps with size $512{\times}512{\times}3$, and is trained with supervision from densely sampled normal maps from the training data.
3.3.2 Visibility Maps
@@
We make the same assumptions to store visibility information as in the case with the UV representation for normals. However, the visibility for each point in 3D requires querying the neural field thousands of times while casting rays from the intersecting point to all the light sources, which is prohibitive during training. Therefore, we pre-compute the visibility UV maps for each pose by sampling multiple camera views and aggregate the results in UV map following Eq. 5. The resulting sparse visibility maps with size $256\times{256}\times{512}$ are then inpainted by the VisibilityNet (as shown in Fig. 2), which takes as input sparse visibility and normal maps, as the visibility highly depends on normals (e.g. the visibility is always zero for the backlight source), and is trained with densely sampled visibility from the training sequence.
3.3.3 Albedo and Roughness Maps
@@
As discussed above, the normal and visibility components can be directly obtained from the density field. However, the albedo and roughness values cannot be explicitly determined in our setup. In intrinsic decomposition, the material properties are commonly optimized via a differentiable renderer. Hence, we optimize the albedo and roughness of the neural actor by leveraging a differentiable renderer (Sec. 3.4) and by imposing priors through our design.
@@
Assuming that the material properties of an actor should not be dependent on the pose, we model albedo and roughness as static UV maps, respectively defined as ${\mathbf{A}\in\mathbb{R}^{1024\times{1024}\times{3}}}$ and ${\mathbf{B}\in\mathbb{R}^{1024\times{1024}}}$. These tensors are jointly optimized for the whole training sequence and regularized with an L1 loss on the spatial gradients, which encourages smoothness in both maps:
@@
where $\nabla_{x}$ and $\nabla_{y}$ represent the spatial gradient operator.
Nonetheless, by assuming a material representation that does not depend on the skeletal pose, we still have to cope with intrinsic mesh registration errors. It is important to mention that, as illustrated in Fig. 2, the human mesh provides guidance to the density field, but also provides the mapping from surface to the UV map. Such registration errors could cause blurry results or artifacts in the final renderings. To cope with this problem, we propose a corrective term to the UV mapping, which is predicted by the UVDeltaNet, as shown in Fig. 2. The UVDeltaNet is an MLP that takes as input the local coordinates $(u,v,h)$ for each point on the ray and the sampled texture features $\psi$, and outputs a corrective UV term $\mathbf{\epsilon}=(u^{\prime},v^{\prime})$.
@@ 3.4 Neural Renderer
The third part of our method is our neural renderer, which takes the disentangled components as normal, visibility, albedo, and rough; integrates the lights from the environment map, and outputs the rendered image.
Based on our volume rendering defined in Sec. 3.2, our neural renderer starts by recasting the ray $\mathbf{r}$ with importance sampling around the estimated depth value $\hat{\mathbf{D}}$. Thus, all the points $\mathbf{r}_{n}$ are most likely close to the surface. For each sampled point in the ray, we obtain the normal direction $\mathbf{n}_{n}$ and the visibility $\bm{\upsilon}_{n}$ by interpolating the normal and visibility maps in UV around the point $(u_{n},v_{n})$. The albedo $\mathbf{a}_{n}$ and roughness $\rho_{n}$ are obtained by interpolating the albedo and roughness maps in UV around the point $(u_{n}+u^{\prime}_{u},v_{n}+v^{\prime}_{n})$.
Before rendering the final image of the neural actor, we consider the light transport equation defined as
$\mathbf{L}_{0}(\mathbf{r}_{n},\bm{\omega}_{0})=\sum_{\bm{\omega}_{i}}\textsc{R}(\mathbf{a}_{n},\rho_{n},\bm{\omega}_{i},\bm{\omega}_{0})\bm{\upsilon}_{n}\mathbf{L}_{i}(\bm{\omega}_{i})(\bm{\omega}_{i}{\mathbf{n}_{n}}),$ (8)
where $\textsc{R}(\cdot)$ is the microfacet BRDF model [42], $\bm{\omega}_{0}$ and $\bm{\omega}_{i}$ are the camera and the light source directions, respectively, and $\mathbf{L}_{i}$ is the incoming light from the environment map. Eq. 8 is defined for a single point $\mathbf{r}_{n}$ on the ray. We integrate the Eq. 8 in our volume rendering as
@@
where $\tilde{\mathbf{I}}$ is the output image of the neural actor.
During training, our relightable model is supervised with the following combined loss:
@@
with $\lambda_{\text{L2}}=100$, $\lambda_{\text{vgg}}=0.01$, and $\lambda_{\text{smooth}}=1$.
@@ 4 Relightable Dynamic Actors Dataset
Figure 4: Samples from our Relightable Dynamic Actors dataset. For each actor, we recorded six sequences under different light conditions, including five indoor sequences and one outdoor. Light environment maps were computed from HDR pictures of a mirror sphere.
Existing datasets for human reenactment and relighting are often monocular [2, 13], contain very short sequences with only a few pose variations, and do not provide environment maps [25, 34, 2, 13]. Hence, we cannot use them for evaluation under realistic scenarios with real ground truth. In addition, we observe that the only existing synthetic dataset for human relighting and pose control [13] suffers from artifacts in the geometry and salt-and-pepper noise in the renderings; see examples in the supplementary material.
With these shortcomings in mind, we collect a new real dataset, which we call Relightable Dynamic Actors, with four identities performing a series of ten different actions. Each subject is recorded six times wearing the same clothes but under different light conditions, including indoor and outdoor, where we used around $100$ and eight cameras, respectively. For each sequence, we also capture HDR light probes, which can be used for training or testing. In total, our dataset contains around $90$k multi-view video frames divided into $24$ multi-view video sequences. See Fig. 4 for samples of our data and the supplement for more details.
@@ 5 Experiments
In this section we provide an experimental evaluation of our method on the Relightable Dynamic Actors dataset and show qualitative and quantitative results. Please see the supplementary video for dynamic results.
@@ 5.1 Evaluation
Metrics. For quantitative evaluation, we report results on the PSNR, SSIM and LPIPS metrics [48], comparing the rendered relight results with the ground-truth images from our dataset. Importantly, our work is the first to be able to report such metrics on a real dataset since no previous dataset in the literature provided a real human recorded under different and known light conditions. Furthermore, since the only method in the literature capable of reposing a human under new light conditions does not provide implementation [13], we defined two baselines for comparisons.
Baselines. Our two baselines are based on our controllable geometry model (Sec. 3.2) and on the obtained normal and visibility, as discussed in Secs. 3.3.1 and 3.3.2. However, instead of using NormalNet and VisibilityNet to obtain these components, we densely sample the neural field and perform morphological inpainting to obtain dense UV maps. The first baseline, referred to as Neural Actor+IR, is an extension of Neural Actor [25] to predict the albedo and rough components instead of RGB color directly from the deformed canonical space $\mathbf{x}_{n}=(x,y,z)$ (see Fig. 2). The second baseline, Neural Actor+IR+Tex, also takes as input the texture features $\psi$ as a conditioning for predicting the albedo and rough values. In both cases, the predictor is a standard NeRF MLP with 8 layers.
Implementation Details. Our model is trained on a three-stage process. First, we train the geometry part (Sec. 3.2) with the Adam [18] optimizer with a learning rate $0.001$ until convergence, which takes about two days on four NVidia A40 GPUs. Then, we pre-train the NormalNet and VisibilityNet with pre-compute normal and visibility UV maps sampled from the geometry model from $40$ views. Both networks are trained in parallel with Adam and the learning rate of $0.01$ for about $12$ hours. Finally, we train our intrinsic decomposition model with the same optimizer until convergence, which takes about one day. The predicted texture maps are obtained with a V2V model [45] trained (in parallel to our model) to translate a sequence of SMPL normal maps into a sequence of texture maps. Please refer to our supplementary materials for architecture details.
@@ 5.2 Results
Figure 5: Qualitative results of our method on novel poses.
Figure 6: Results from our intrinsic decomposition.
Figure 7: Our method models global illumination and cast shadows, as can be observed from our results under “sunrise” light.
@@
Table 2: Quantitative results from our method on real data, subject S1, considering indoor and outdoor scenes. Our method consistently improves the baselines, especially on the outdoor sequence, where the gap between training and test light conditions is more relevant.
Qualitative results. In Fig. 5, we show qualitative results of our method. Our results are all from a trained model driven by a new skeletal pose, new camera viewpoint, and new environment map not seeing during training. Note how our approach captures realistic appearance and prominent cloth wrinkles, which can be easily observed under different light conditions. Importantly, notice how our results look realistic under novel light conditions.
In Fig. 6, we show results from our intrinsic decomposition. Note how our normal representation captures well the geometry details, and how our visibility captures shadows. This is important to obtain an albedo that does not embed light effects, as can be observed from our results. Another important and unique characteristic of our method is the ability to model self-shadows. In Fig. 7, we show results from two subjects, where a “sunrise” environment map was used to produce cast shadows.
Comparisons. We evaluated our method on the test sequences from our dataset. Specifically, we compare with two baselines in the indoor sequences (T2-T5) and outdoor sequence (T6) of subject S1 in every $10$th frame, which sums up to $4000$ evaluation frames. In Tab. 2 we show, for the first time, quantitative results of real human relighting under new poses and new light conditions. Our approach consistently improves upon the two baselines.
Ablation. We show in Tab. 3 an ablation with different components of our method. Compared to our baseline (Neural Actor+IR), learning the albedo and roughness as UV maps significantly improves the performance on relighting and reposing. This is because, in the baselines, the albedo and roughness are under-constrained and can possibly embed shading effects, overfiting to the training data. The NormalNet and VisiblityNet provide additional capacity to the model to learn how to inpaint the sampled UV maps. Note that without these two models our results are computed using morphological region growth in the UV maps. Finally, the UVDeltaNet provides more flexibility to the albedo and roughness representation, accounting for fine details. This can be observed in the improvement in the LPIPS metric in the last row of Tab. 3. Nonetheless, this also means more capacity to overfit the training sequences, which translates into a small decrease in the PSNR metric.
@@
Table 3: Ablation study with different components of our method.

### v2

@@ 3.2 Neural Intrinsic Decomposition
At the core of our method, we perform intrinsic decomposition to obtain the normal, visibility, albedo, and roughness components, required for rendering the neural actor under new light conditions. Specifically, for each point in each ray cast into the scene we want to obtain the normal direction $\mathbf{n}\in\mathbb{R}^{3}$, the visibility coefficients $\bm{\upsilon}\in\mathbb{R}^{512}$ respective to each light source in the environment map, the RGB albedo $\mathbf{a}\in\mathbb{R}^{3}$, and the scalar roughness $\rho$. The normal and visibility values can be computed from the geometry model. However, computing the visibility of each point in 3D with respect to the light sources is computationally intractable for a dynamic and long sequence, since for every point in 3D and for every frame, several rays have to be cast towards each pixel of the environment map. Therefore, we propose a different approach.
@@
We notice that the intrinsic decomposition is only relevant for points close to the surface. Since the coarse human mesh provides a mapping from the mesh surface to UV space and the implicit human model provides a density field that models the geometry of the actor, we propose an aggregation strategy that samples normal and visibility from a given point of view (therefore, only a few points are sampled in 3D) and aggregates the obtained values in a per-frame UV map. Once we obtain the normal and visibility in UV space for a given pose, we can simply query the values by projecting the sampled 3D point $\mathbf{r}_{n}$ into the UV space, preventing our method from the expensive computation of normal and visibility at every possible 3D coordinate over multiple frames. We next explain this approach in more detail.
3.2.1 Normal Maps.
@@
For each point in a ray, we want to obtain the normal value that will be required for solving the rendering equation during relighting (will be introduced in Sec. 3.3). This value could be obtained by computing the gradient from the neural field [63]. However, this process requires back-propagating gradients throughout the entire network and often results in noisy normals. Instead, we explicitly compute the normal vector close to the visible surface by deriving the unbiased depth $\hat{\mathbf{D}}$ [54]:
$\hat{\mathbf{D}}(\mathbf{r})=\frac{\sum_{n=1}^{N}{w_{n}\delta_{n}}}{\sum_{n=1}^{N}{w_{n}}},\quad\text{with}\;w_{n}=T_{n}\left(1-\text{exp}\left(-\sigma_{n}{\bar{\delta}_{n}}\right)\right),$ (1)
where $T_{n}=\text{exp}\left(-\sum_{j=1}^{n-1}\sigma_{j}{\bar{\delta}_{j}}\right)$, $w_{n}$ is the density weight from volumetric rendering [39], and $\sigma_{n}$ is the density predicted by the NeRF model at point $\mathbf{r}_{n}$.
@@
Figure 3: The unbiased depth from Eq. 1 is projected into 3D coordinates, where we approximate the tangent surface at the intersecting point and obtain the normal vector. Normals are sampled from multiple viewpoints and aggregated in the UV map.
@@
Then, as illustrated in Fig. 3, we convert the depth values relative to the camera to 3D coordinates and obtain the normal direction from the tangent surface computed at the intersection point, defined as the maximum value of the density weight $w_{n}$. This process can be performed for multiple viewpoints to cover the surface of the human.
@@
Once the normal values are obtained in 3D, we project them to the UV map and aggregate them with the obtained normals from different views. In this process, the $(u^{*},v^{*})$ coordinate where the obtained normal will be projected onto the UV map is defined by
$(u^{*},v^{*})=\Psi(\mathbf{r}_{n}^{*}),\quad\text{with}\;\mathbf{r}_{n}^{*}=\operatorname*{arg\,max}_{\mathbf{r}_{n}}w_{n},$ (2)
@@
where $\Psi$ is a function that projects a 3D point to the closest surface point in the human mesh. Note that $w_{n}$ depends on the sampled points $\mathbf{r}_{n}$ (from $T_{n}$ in Eq. 1) and $(u^{*},v^{*})$ corresponds to the $(u_{n},v_{n})$ value obtained from the point $\mathbf{r}_{n}$ on the ray that has the largest density weight $w_{n}$. This process ensures that the normal map obtained aggregates the normals most likely close to the surface.
@@
In this process, the resulting UV map can be sparse, which is undesirable for free-viewpoint rendering. Thus, the NormalNet is used to inpaint the missing normals in UV space. This lightweight CNN is implemented with $8$ layers of partial convolutions [31], takes as input sparse normal maps with size $512{\times}512{\times}3$, and predicts a dense normal map (see Fig. 2). During training, NormalNet is supervised with densely sampled normal maps from the training data.
3.2.2 Visibility Maps.
@@
The visibility information is necessary to model pose-dependent shadows in the neural avatar, without which it is not possible to produce realistic relighting results. A naive approach for obtaining the visibility for each point in 3D would require querying the neural field thousands of times while casting rays from the intersecting point to all the light sources, which is prohibitive during training. In contrast, in our approach we efficiently encode visibility of each point in 3D close to the surface by applying an approximate mapping to UV space. Specifically, we pre-compute the visibility maps for each pose by sampling multiple camera views and aggregate the results in UV map following Eq. 2. The resulting visibility maps with size $256\times{256}\times{512}$ encode the visibility information w.r.t. each light source in the environment map (represented by $512$ individual lights). Similarly to the computation of normal maps, this process results in sparse visibility maps, which are inpainted by the VisibilityNet (as shown in Fig. 2). This shallow CNN takes as input sparse visibility and normal maps, as the visibility highly depends on normals (e.g. the visibility is always zero for the backlight source), and is trained with densely sampled visibility from the training sequence. See the suplement for archtecture details.
3.2.3 Albedo and Roughness Maps.
@@
As discussed above, the normal and visibility components can be computed from the density field by using UV maps as proxies. However, the albedo and roughness properties cannot be directly obtained from the pre-trained Neural Actor model, since it predicts radiance which is the product of the rendering equation. Instead, we want to learn a decomposed representation. In intrinsic decomposition, the material properties are commonly optimized via differentiable rendering. Hence, we optimize the albedo and roughness of the neural actor by leveraging a differentiable renderer (Sec. 3.3) and by imposing priors through our design.
@@
The material properties of an actor should not depend on the pose and should be efficiently encoded to enable fast computations and editing. Therefore, we model albedo and roughness as static UV maps, respectively defined as ${\mathbf{A}\in\mathbb{R}^{1024\times{1024}\times{3}}}$ and ${\mathbf{B}\in\mathbb{R}^{1024\times{1024}}}$. These tensors are jointly optimized for the whole training sequence and regularized with an L1 loss on the spatial gradients, which encourages smoothness in both maps:
@@
where $\nabla_{x}$ and $\nabla_{y}$ represent the spatial gradient operators.
Nonetheless, by assuming a material representation that does not depend on the skeletal pose, we still have to cope with intrinsic mesh registration errors. Such registration errors could cause blurry results or artifacts in the final renderings. To cope with this problem, we propose a corrective term to the UV mapping, which is predicted by the UVDeltaNet, as shown in Fig. 2. The UVDeltaNet is an MLP that takes as input the local coordinates $(u,v,h)$ for each point on the ray and the sampled texture features $\psi$, and outputs a corrective term $\mathbf{\epsilon}=(u^{\prime},v^{\prime})$. The texture features are necessary to provide additional information, since the local coordinates alone cannot account for deviations in the mesh registration.
@@ 3.3 Neural Renderer
The third part of our method is our neural renderer, which takes the disentangled components as normal, visibility, albedo, and roughness; integrates the lights from the environment map, and outputs the rendered image.
Our neural renderer starts by recasting the ray $\mathbf{r}$ with importance sampling around the estimated depth value $\hat{\mathbf{D}}$. Thus, all the points $\mathbf{r}_{n}$ are most likely close to the surface. For each sampled point in the ray, we obtain the normal direction $\mathbf{n}_{n}$ and the visibility values $\bm{\upsilon}_{n}$ by interpolating the normal and visibility maps in UV around the point $(u_{n},v_{n})=\Psi(\mathbf{r}_{n})$. The albedo $\mathbf{a}_{n}$ and roughness $\rho_{n}$ are obtained by interpolating the albedo and roughness maps in UV around the point ${(u_{n}+u^{\prime}_{u},v_{n}+v^{\prime}_{n})}$, i.e., following the deformation predicted by UVDeltaNet.
Before rendering the final image of the posed neural actor, we consider the light transport equation defined as
$\mathbf{L}_{0}(\mathbf{r}_{n},\bm{\omega}_{0})=\sum_{i}\textsc{R}(\mathbf{a}_{n},\rho_{n},\bm{\omega}_{i},\bm{\omega}_{0})\bm{\upsilon}_{n,i}\mathbf{L}(\bm{\omega}_{i})(\bm{\omega}_{i}{\mathbf{n}_{n}}),$ (5)
[…]
@@
where $\tilde{\mathbf{I}}$ is the output image of the relighted neural actor.
During training, our relightable model is supervised with the following loss:
@@
where $\mathcal{L}_{\text{L2}}$ and $\mathcal{L}_{\text{vgg}}$ are the L2 and the perceptual losses [26] between $\tilde{\mathbf{I}}(\mathbf{r})$ and the reference image. The hyperparameters are defined as $\lambda_{\text{L2}}=100$, $\lambda_{\text{vgg}}=0.01$, and $\lambda_{\text{smooth}}=0.1$ in all our experiments.
@@ 4 Relightable Dynamic Actors Dataset
Figure 4: Relightable Dynamic Actors dataset. For each actor, we recorded six sequences under different light conditions, including five indoor sequences and one outdoor. Environment maps were obtained from HDR pictures of a mirror sphere.
Existing datasets for human reenactment and relighting are often monocular [2, 17], contain very short sequences with only a few pose variations, and do not provide environment maps [32, 42, 2, 17]. Hence, we cannot use them for evaluation under realistic scenarios with real ground truth. In addition, we observe that the only existing synthetic dataset for human relighting and pose control [17] suffers from artifacts in the geometry and salt-and-pepper noise in the renderings.
With these shortcomings in mind, we collect a new real dataset, which we call Relightable Dynamic Actors, with four identities performing a series of ten different actions (Fig. 4). Each subject is recorded six times wearing the same clothes but under different light conditions, including indoor and outdoor, where we used around $100$ and eight cameras, respectively. For each sequence, we also capture HDR light probes, which can be used for training or testing. In total, our dataset contains around $90$k multi-view video frames divided into $24$ multi-view video sequences. We refer to the supplemental document for more details.
@@ 5 Experiments
In this section, we provide an experimental evaluation of our method on the Relightable Dynamic Actors dataset and show qualitative and quantitative results. Please refer to the supplementary video for dynamic results on long sequences.
@@ 5.1 Evaluation
Metrics. For quantitative evaluation, we report results on the PSNR, SSIM and LPIPS metrics [62], comparing the rendered relight results with the ground-truth images from our dataset. Importantly, our work is the first to be able to report such metrics on a real dataset since no previous dataset in the literature provided a real human recorded under different and known light conditions.
Baselines. Since the only method in the literature capable of reposing a human under new light conditions does not provide implementation [17], and Relighting4D [6] is not able to perform inference for new poses, we defined two baselines for a fair comparison. Our baselines are derived from Neural Actor [32] and adapted to perform intrinsic decomposition based on the obtained normal and visibility information, as discussed in Sec. 3.2.1. However, instead of using NormalNet and VisibilityNet to obtain these components, we densely sample the neural field and perform morphological inpainting to obtain dense UV maps. The first baseline, referred to as Neural Actor+IR, predicts the albedo and roughness components instead of RGB color directly from the deformed canonical space $\mathbf{x}_{n}=(x,y,z)$ (see Fig. 2). The second baseline, Neural Actor+IR+Tex, takes as input the texture features $\psi$ as a conditioning for predicting the albedo and roughness values. In both cases, the predictor is a NeRF MLP with 8 layers.
Implementation Details. Our model is trained in a three-stage process. First, we train the geometry component (Sec. 3.1) with the Adam [23] optimizer with a learning rate of $0.001$ until convergence, which takes about two days on four NVidia A40 GPUs. Then, we pre-train the NormalNet and VisibilityNet with pre-computed normal and visibility UV maps sampled from the geometry model from $40$ different views. Both networks are trained in parallel using Adam with a learning rate of $0.01$ for about $12$ hours. Finally, we train our intrinsic decomposition model with the same optimizer until convergence, which takes about one day. Since the obtained environment maps are up to factor of scale [8], we apply global color correction during training and evaluation. The predicted texture maps are obtained with a V2V model [56] trained (in parallel to our model) to translate a sequence of SMPL normal maps into a sequence of texture maps. Please refer to our supplementary materials for architecture details.
@@ 5.2 Results
Figure 5: Qualitative results of our method on novel poses. Our approach generalizes to both indoor and outdoor illumination, and generates realistic renderings for arbitrary environment maps with detailed shadows and pose-dependent deformations.
Qualitative Results. In Fig. 5, we show qualitative results of our method. Our results are all from a trained model driven by a new skeletal pose, new camera viewpoint, and new environment map not seen during training. Note how our approach produces realistic appearance and prominent cloth wrinkles, which can be easily observed under different light conditions. Importantly, notice how our results look realistic under novel lights, even though our model is trained on a single sequence under a static light condition.
Figure 6: Results from our intrinsic decomposition. Our method provides detailed normals and visibility (averaged for visualization), and plausible albedo and roughness predictions that enable realistic relighting.
Table 1: Quantitative results from our method on real data, subject S1, considering indoor and outdoor scenes. Our method consistently improves the baselines, especially in outdoor scenarios, where the gap between training and test lights is more pronounced.
@@
In Fig. 6, we show results from our intrinsic decomposition. Note that our normal representation nicely captures the geometry details and the visibility. This is important to obtain an albedo that does not contain baked-in lighting effects, e.g. shadows, as can be observed from our results.
Figure 7: Our method models pose-dependent effects and cast shadows, as can be observed from our results under the “sunrise” light.
Another important and unique characteristic of our method is the ability to model self-shadows and cloth wrinkles. In Fig. 7, we show results from two subjects, where a “sunrise” light probe was used to produce cast shadows. Note how pose-dependent effects are consistent with the mostly directional illumination.
Comparisons. We evaluated our method on the test sequences from our dataset. Specifically, we compare against two baselines in the indoor sequences (T2-T5) and outdoor sequence (T6) of subject S1 in every $10$th frame, which sums up to $4000$ evaluation frames. In Tab. 1 we show, for the first time, quantitative results of real human relighting under new poses and new light conditions. Our approach consistently improves upon the two baselines.
Ablation. In Tab. 2, we show an ablation with different components of our method. Compared to our baseline (Neural Actor+IR), learning the albedo and roughness as UV maps guided by the neural field significantly improves the performance on relighting and reposing. This is because, in the baselines, the albedo and roughness are under-constrained and can possibly embed shading effects, overfitting to the training data. The NormalNet and VisiblityNet provide additional capacity to the model to learn how to inpaint the sampled UV maps. Note that without these two models our results are computed using morphological region growth in the UV maps. Finally, the UVDeltaNet provides more flexibility to the albedo and roughness representation, accounting for fine details. This can be observed in the improvement in the LPIPS metric in the last row of Tab. 2. Nonetheless, this also means more capacity to overfit the training sequences, which translates into a small decrease in the PSNR metric.
Table 2: Ablation study with different components of our method. The proposed intrinsic decomposition with static albedo and roughness maps significantly improves over the baseline, and the proposed NormalNet and VisibilityNet further improve the results while enabling free-viewpoint rendering. UVDeltaNet provides more realistic results (LPIPS metric) with a small sacrifice in the PSNR.